solve x^4-8x^2+7=0
Answer by venugopalramana(3286)

solve x^4-8x^2+7=0
X^4-X^2-7X^2+7=0
X^2(X^2-1)-7(X^2-1)=0
(X^2-1)(X^2-7)=0
(X-1)(X+1)(X-SQRT7)(X+SQRT7)=0
X=1..........OR
X=-1....OR
X=SQRT7........OR
X=-SQRT7